Project

General

Profile

SamWiseMinutes3 20 14

Gavin, Kanika, Craig, Nate, Ryan, Dominick, Jan, Eric,

Action Items:
——————
—> Get status from Eric
—> New tag (or patches) for Chris with photon transport update?
— Ruth and Nate (mixing) have something too…
—> Get Jim new mixer files:
— Requires the patch – Mixer, photon transport, Need commit from Ruth.
—> Under pressure from Rick for partition 2.
—> Pressure from Gavin for reconstruction for neutrino search

FTS — transferring 600-1000 files per hour.
—> No big slow downs lately.
—> Lots waiting, but a lot pushed through lately. (10,000 files pending)

How are we going to process data? Do we want reco files on BlueArc?
—> Lots of collaboration will want to use them.
—> Will the tools to process reco files be accessible to the collaboration.
—> Some CAF files will be produced directly off of RECO.

Production status:
—> Eric’s files are finished – we don’t know the success rate.
—> We got banned from OSC. Hi memory usage – maybe it wasn’t us? Running again now.
—> Will produce ND cosmics on amazon cloud
—> Some issues with reco came up from the reco validation.

New issue with backtracker where some objects don’t have a parent.
—> Not sure what is going on here?
—> Reco new tag “up in the air” until we can understand

Mixer: (Nate and Jim)
—> Bad mixing files due to process name in fcl files
—> Fix is committed
—> Update to mixer file to pick up the MC-flux information
(backport or patch required to produce new test file for Jim)
—> Can run jobs on the grid now.

Multi art jobs per grid job: (Dominick)
—> Got them up and running yesterday
—> Full scale usage this morning – working great. 1.5 hours for 5 jobs. Copy back is slow using BESTMAN.
(may not need BESTMAN when running at Fermilab)
—> Do we have a way to monitor BESTMAN?
—> Most data since late February processed within a day.

Status of data:
- PCHits are moving along.
- Raw2root also going well.
- Will need to hit the switch on reco pretty soon for the neutrino search.

Action Items from last time:
———————————————
— Currently files are still sitting on Blue Arc.
—> implemented at FTS (only PID and CAF.root go to BlueArc only)
—> What about the drop boxes?

— Ready to run Rock Files and FD GENIE (4 flavors)
- Need drop box and fhcl files
- Some will be run on OSC, HARVARD, SMU, …
- Last time, Fermilab took all the jobs…
—> Submitted, but need an update from Eric

— Reco next, but we need new calibration constants of the MC
— Status of calibration:
- No SAM or batch processing problems!
- Physics problems